# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ 1 Introduction
Our theoretical results highlight the practical relevance of cryptographic tools—particularly zero-knowledge proofs (ZKPs) [18, 40]—in the context of distributed learning. ZKPs can prevent Byzantine failures by verifying that each worker’s input data lies within an appropriate range and that their model updates remain consistent with their initially committed dataset, all without revealing the data itself [35, 1, 37]. However, they do not guard against data poisoning, which stems from malicious or corrupted local data. Therefore, cryptographic safeguards must be paired with robust learning algorithms that can tolerate compromised training data to ensure end-to-end resilience.
2 Preliminaries
2.1 Problem setting
We consider a distributed setup with a central server and $n$ workers, where up to $f$ (with $f<\frac{n}{2}$) may be subject to either Byzantine failure or data poisoning attacks, as defined below. We refer to such workers as misbehaving, unless we need to distinguish between the two attack types. Their identities are unknown to the server. While the actual number of misbehaving workers may be less than $f$, we assume the worst-case scenario where exactly $f$ workers are misbehaving. The remaining honest workers form the set $\mathcal{H}\subseteq[n]\vcentcolon=\{1,\ldots,n\}$, with $|\mathcal{H}|=n-f$. Each honest worker $i\in\mathcal{H}$ holds a local dataset $\mathcal{D}_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}$ composed of $m$ i.i.d. data points (or samples) from an input space $\mathcal{Z}$ drawn from a distribution $p_{i}$. We denote $\mathcal{S}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Given a parameter vector $\theta\in\Theta\subset\mathbb{R}^{d}$ representing the model, a data point $z\in\mathcal{Z}$ incurs a differentiable loss defined by a real-valued function $\ell(\theta;z)$. The goal is to minimize the population risk over the honest workers:
As we only have access to a finite number of samples from each distribution, the above objective can only be solved approximately by minimizing the empirical risk over the honest workers:
$\widehat{R}_{\mathcal{H}}(\theta)=\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\widehat{R}_{i}(\theta),\quad\text{with}\quad\widehat{R}_{i}(\theta)=\frac{1}{m}\sum_{z\in\mathcal{D}_{i}}\ell(\theta,z).$
The population risk $R_{\mathcal{H}}(\theta)$ of a distributed learning algorithm’s output $\mathcal{A(\mathcal{S})}$ can be decomposed into the generalization and optimization errors (see, e.g. [9, 23]), i.e.,
$\mathbb{E}\big{[}R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big{]}\leq\mathbb{E}\big{[}R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big{]}+\mathbb{E}\big{[}\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}\widehat{R}_{\mathcal{H}}(\theta)\big{]}$ (1)
Prior work primarily only considers the optimization error, i.e., the second term on the right-hand side. In contrast, we are interested in minimizing the honest population risk under Byzantine attacks. Formally, we define the notion of robustness as follows. We denote the Euclidean norm by $\|\cdot\|_{2}$.
[…]
A distributed algorithm $\mathcal{A}$ is said to be $(f,\rho,\text{statistical})$-resilient if it enables the server to output a parameter vector $\hat{\theta}$ such that $\mathbb{E}_{\mathcal{A}}\big{[}R_{\mathcal{H}}(\hat{\theta})-\inf_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big{]}\leq\rho$. In the case of nonconvex population risk, we aim for $\mathbb{E}_{\mathcal{A}}\|\nabla R_{\mathcal{H}}(\hat{\theta})\|_{2}^{2}\leq\rho$. We say $(f,\rho,\text{empirical})$-resilient when we instead consider empirical risk.
Note that the (standard) notion of $(f,\rho,\text{empirical})$-resilience—see, e.g., [30]—is impossible in general (for any $\rho$) when $f\geq n/2$, justifying our assumption $f<n/2$.
Threat models. We aim to analyze this resilience property under two standard threat models widely studied in the literature [14, 38].
Byzantine failure attacks. In this threat model, misbehaving workers can act arbitrarily: they may deviate from the prescribed algorithm and send adversarial updates to the server. These Byzantine workers are omniscient, with access to all information exchanged between honest workers and the server. This model follows the classical notion of Byzantine failures in distributed systems introduced in the seminal work of [27].
Data poisoning attacks. In this restricted threat model, misbehaving workers follow the prescribed algorithm but may arbitrarily corrupt their local training data before execution. Specifically, for each misbehaving worker $i\notin\mathcal{H}$, a local dataset $\mathcal{D}_{i}$ is sampled from a distribution $q_{i}$ over $\mathcal{Z}$, where $q_{i}$ is adversarially chosen with full knowledge of the honest workers’ data distributions $\{p_{i},~{}i\in\mathcal{H}\}$.
2.2 Background on robust distributed optimization
Robust distributed optimization algorithms aim to achieve $(f,\rho,\text{empirical})$-resilience. They are typically adaptations of standard first-order iterative optimization algorithms like gradient descent ($\mathrm{GD}$) or stochastic gradient descent ($\mathrm{SGD}$). These algorithms are made robust by replacing the server-side averaging operator by a robust aggregation rule $F:(\mathbb{R}^{d})^{n}\to\mathbb{R}^{d}$ [20]. Formally, given a learning rate $\gamma$, the parameter $\theta_{t}$ at iteration $t\in\{0,\ldots,T-1\}$ is updated as follows:
where $g^{(i)}_{t}$ denotes the update sent by worker $i$ at iteration $t$. For an honest worker $i\in\mathcal{H}$, we have $g^{(i)}_{t}=\nabla\widehat{R}_{i}(\theta_{t})$ under $\mathrm{GD}$, or $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i)}_{t})$ under $\mathrm{SGD}$, with $z^{(i)}_{t}$ uniformly sampled from the local dataset $\mathcal{D}_{i}$. For a misbehaving worker $i\notin\mathcal{H}$, $g^{(i)}_{t}$ is either an arbitrary vector in $\mathbb{R}^{d}$ (Byzantine failure) or a gradient computed on corrupted data (poisoning attack).
Recent advances in robust distributed optimization have identified key properties that a robust aggregation rule must satisfy to ensure $(f,\rho,\text{empirical})$-resilience [4, 5]. The following definition encompasses a wide range of aggregation schemes and has been shown to yield tight resilience guarantees across a variety of practical distributed learning settings.
Definition 2.2 ($(f,\kappa,\|\cdot\|_{op})$-robustness).
Let $\|\cdot\|_{op}$ denote an operator norm for p.s.d. matrices. Let $n\geq 1$, $20\leq f<n/$ 2 and $\kappa\geq 0$. An aggregation rule $F$ is $(f,\kappa,\|\cdot\|_{op})$-robust if for any $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, any set $S\subset[n]$ of size $n-f$, with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g_{i}$ and $\Sigma_{S}=\frac{1}{|S|}\sum_{i\in S}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$,
$\|F(g_{1},\ldots,g_{n})-\overline{g}_{S}\|_{2}^{2}\leq\kappa\|\Sigma_{S}\|_{op}.$
Depending on the context, we either use the nuclear norm (i.e., the trace $\operatorname{Tr}$) or the spectral norm $\|\cdot\|_{sp}$ (i.e., the largest eigenvalue) for $\|\cdot\|_{op}$. $f$ and $\kappa$ are referred to as the robustness parameter and robustness coefficient of $F$, respectively.
An aggregation rule $F$ that is $(f,\mathcal{O}(f/n),\|\cdot\|_{sp})$-robust—such as smallest maximum eigenvalue averaging ($\mathrm{SMEA}$, which averages the $n-f$ gradients that are most directionally consistent, discarding outliers based on covariance; see Definition C.1)—achieves optimal $(f,\rho,\text{empirical})$-resilience. Specifically, iterative methods using such $F$ attain an optimization error that matches the information-theoretic lower bound for first-order optimization under $f$ Byzantine workers [5]. When the empirical covariance matrix of honest workers’ gradients has uniformly bounded trace over $\Theta$, tight optimization error guarantees can also be obtained under the weaker $(f,\mathcal{O}(f/n),\operatorname{Tr})$-robust property. This weaker condition can be satisfied by computationally cheaper rules, such as coordinate-wise trimmed mean ($\mathrm{CWTM}$, which removes the largest and smallest $f$ values in each coordinate and averages the remaining entries; see Definition C.2[4]. As noted in our comparison to related work (Section 5), prior analyses focus on optimization error, failing to account for discrepancies in generalization performance under different threat models. We address this limitation by studying how robust aggregation impacts the stability of Byzantine-resilient distributed learning algorithms.
2.3 Stability under misbehaving workers
The main objective of this work is to analyze the generalization error (eq. 1) of robust distributed optimization algorithms. We do so through the lens of uniform algorithmic stability, which measures how sensitive a learning algorithm is to changes in its training data. This approach originates in early works [41, 12, 33] and was later popularized in [10, 36, 23]. We revisit this framework to study generalization in robust distributed learning, where the change is only in the honest workers’ data.
Definition 2.3 (Stability with misbehaving workers).
Consider $n$ workers, among which $f$ are misbehaving, and $n-f$ are honest with local datasets of size $m$. A distributed randomized algorithm $\mathcal{A}$ is said to be $\varepsilon$-uniformly stable if for all pairs of honest datasets $\mathcal{S},\mathcal{S}^{\prime}\in\mathcal{Z}^{(n-f)m}$ that differ in at most one sample (referred to as neighboring datasets), we have $\sup_{z\in\mathcal{Z}}\mathbb{E}[\ell(\mathcal{A}(\mathcal{S});z)-\ell(\mathcal{A}(\mathcal{S}^{\prime});z)]\leq\varepsilon$, where the expectation is over the randomness of $\mathcal{A}$.
This property leads to the following bound on the generalization error, established in Appendix A.

### v2

@@ 6 CONCLUSION
Our theoretical results highlight the relevance of cryptographic tools—particularly zero-knowledge proofs (ZKPs) (Goldwasser et al., 1989; Thaler, 2022)—for distributed learning. ZKPs can prevent Byzantine failures by verifying that each worker’s input data lies within an appropriate range and that their model updates remain consistent with their initially committed dataset, all without revealing the data itself (Sabater et al., 2022; Abbaszadeh et al., 2024; Shamsabadi et al., 2024). However, they do not guard against data poisoning, which stems from malicious or corrupted local data. Therefore, cryptographic safeguards must be paired with robust learning algorithms that can tolerate compromised training data to ensure end-to-end resilience.
Future work.
Promising directions include: (i) Identify a robust aggregation rule preserving the loss function’s co-coercivity, or prove this is impossible; (ii) Go beyond worst-case scenarios with data-dependent refinements like those based on on-average stability (Lei and Ying, 2020; Sun et al., 2024; Ye et al., 2025); (iii) Extend results to the locally-poisonous setting (Farhadkhani et al., 2024b); (iv) Investigate whether the trilemma between robustness, privacy and optimization error shown in recent work (Allouah et al., 2023b) also holds for generalization error; (v) Quantify the stability of other first-order algorithms, including momentum and communication-efficient methods (see Appendix H for a discussion).
Acknowledgements
Supplementary Materials: Byzantine Failures Harm the Generalization of Robust Distributed Learning Algorithms More Than Data Poisoning
The supplementary materials are organized as follows.
  • Appendix A contains deferred proofs for Section 2, focusing on the link between uniform stability and generalization bounds under Byzantine failures and data poisoning.
  • Appendix B summarizes the regularity assumptions and includes key expansivity results.
  • Appendix C provides detailed proofs of Theorems 3.1 and 3.2 presented in Section 3.1.
  • Appendix D provides detailed proofs of Theorems 3.3 and 3.4 presented in Section 3.2.
  • Appendix E specifies details supporting Section 4, including generalization error computations (Section E.1) and numerical experiments (Section E.2).
  • Appendix F gives example of results with refined bounded heterogeneity and bounded variance assumptions instead of the more general bounded gradients assumption.
  • Appendix G states high-probability excess risk bounds for robust distributed learning under smooth and strongly convex loss functions.
  • Appendix H discusses additional important points not included in the main text.
Appendix A STABILITY IMPLIES GENERALIZATION UNDER BYZANTINE FAILURES OR DATA POISONING
In this section, we focus on the formal link between uniform stability and generalization under Byzantine failures and data poisoning. The proof follows classic arguments from Bousquet and Elisseeff (2002).